OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. i, 2 - √3

OpenStudy (anonymous):

Expand \[ (x-i) (x+i) \left(x-\left(2-\sqrt{3}\right)\right) \left(x-\left(\sqrt{3}+2\right)\right) \]

OpenStudy (anonymous):

would that be the answer?

OpenStudy (anonymous):

That is the answer \[ (x-i) (x+i) \left(x-\left(2-\sqrt{3}\right)\right) \left(x-\left(\sqrt{3}+2\right)\right)=\\ \left(x^2+1\right) \left(x^2-4 x+1\right)=\\x^4-4 x^3+2 x^2-4 x+1 \]

OpenStudy (anonymous):

If a + ib is a root, then a - ib is also a root. If 2- \sqrt(3) is arootthen 2-\sqrt(3) is a root.

OpenStudy (anonymous):

thank you so much would you know how to do it when it gives 1, 4, 1 + √2

OpenStudy (anonymous):

\[(x-1)(x-4)( x -(1 +\sqrt 2)) ( x -(1-\sqrt 2)) \]

OpenStudy (anonymous):

thank you so very much

OpenStudy (anonymous):

yw